### v3

@@ 4.2 Transforming an OEN into datalog
The transformation takes an OEN and an initial marking $m$. Let $t$ denote any transition. Input places of transitions are represented by hypernodes $p^{*}$. A datalog program $\mathcal{P}$ is written with predicate symbols $t$, $r_{1}$ and $r_{2}$ and variables X, Y and Z:
  • write $token(m,p^{*})$ if the marking $m$ includes all members of a hypernode $p^{*}$
@@
  • write $t(p_{1}^{*},p_{2}^{*})$ if $t$ connects two hypernodes $p_{1}^{*}$ and $p_{2}^{*}$
  • write $t(p^{*},q)$ if $t$ connects a hypernode $p^{*}$ and a place $q$
@@
The transformed program is linear and immediately recursive since it contains only one recursive clause and the recursive predicate appears only once in the head of this clause. We illustrate how reachable places from an initial marking can be computed by evaluating the transformed program.
@@
Let the OEN in Figure 2 represent a metabolic pathway. Transitions $t_{1}$, $t_{2}$ and $t_{3}$ are described by $reaction\_1$, $reaction\_2$ and $reaction\_3$ clauses. We create two hypernodes denoted by $c1\_c2$ and $c3\_c4$. These transitions connect places $\{c_{3}$, $c_{4}$, $c_{5}\}$ which are represented by constants $\{c3$, $c4$, $c5\}$. For an initial marking $m_{1}=\{c_{1},c_{2}\}$, we can construct a datalog program $\mathcal{P}_{1}$:
$\mathcal{P}_{1}:\left\{\begin{aligned} token(m1,c1\_c2).&\\ token(m1,c1).&\\ token(m1,c2).&\\ reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ reaction(X,Y)\leftarrow token(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
[Uncaptioned image][Uncaptioned image]
The right-hand-side figures are the OEN from from Figure 2 and the hypergraph represented by $\mathcal{P}_{1}$. The hypernode $\{c_{3},c_{4}\}$ is reachable from $\{c_{1},c_{2}\}$ via the transition $t_{1}$. The place $c_{5}$ is reachable in two steps via transitions $t_{1}$ and $t_{2}$. We can compute these by evaluating groundings of $pathway(m1,Y)$ in the Herbrand model $M(\mathcal{P}_{1})$ which gives us $pathway(m1,c2)$, $pathway(m1,c3\_c4)$ and $pathway(m1,c5)$.
@@
It follows from Equation 1 that all tokens at the places in the pre-set of a transition are consumed and places in the post-set of a transition are assigned tokens.
@@
Let $(P,T,F)$ be an arbitrary OEN. Using Proposition 1, we show $\mathcal{T}_{F}$ over the lattice $(\mathcal{T}_{F},\subseteq)$ has a fixpoint. We refer to Tarski’s fixpoint theorem tarski_fixpoint , which states a monotonic operator on a complete lattice has a least fixpoint. We prove that $\mathcal{T}_{F}$ is a monotonic operator.
@@
$\mathcal{T}_{F}$ is a monotonic function. According to Tarski’s fixpoint theorem tarski_fixpoint , the function $\mathcal{T}_{F}$ has a least fixpoint $\mathcal{T}_{F}{\uparrow^{\alpha}}$ for some $\alpha\geq 0$. Since $\mathcal{T}_{F}$ is also a compact operator, if $p\in\mathcal{T}_{F}(\mathcal{T}_{F}{\uparrow^{\alpha}})$ then $p\in\mathcal{T}_{F}(\mathcal{T}_{F}{\uparrow^{n}})$ for some $n\geq 0$ and $\mathcal{T}_{F}{\uparrow^{n}}\subseteq\mathcal{T}_{F}{\uparrow^{\alpha}}$. The union of subsets of $\mathcal{T}_{F}{\uparrow^{\alpha}}$ is $\cup_{n\geq 0}\mathcal{T}_{F}{\uparrow^{n}}$, which covers every member of the least fixpoint. Therefore, $\mathcal{T}_{F}{\uparrow^{\alpha}}=\cup_{n\geq 0}\mathcal{T}_{F}{\uparrow^{n}}$.
@@
Prove the forward implication. Assume $\mathcal{P}\models r(c_{1},c_{2})$. Constant symbols $c_{1}$ and $c_{2}$ represent place nodes in the OEN. We know $r(c_{1},c_{2})$ is in the least model. Based on the definition of the Immediate Consequence Operator van_emden_semantics_1976 , $r(c_{1},c_{2})\leftarrow\,B_{1},...,B_{i}$ ($i\geq 0$) is a clause in $\mathcal{P}$ and $\{B_{1},...,B_{i}\}\subseteq\mathcal{T_{P}}{\uparrow^{\omega}}$. From Definition 1, we know that $B_{1},...,B_{i}$ are unit ground clauses describing the pre-sets and post-sets of transitions. There exists $n\geq 0$ such that $c_{1}\in\mathcal{T}_{F}{\uparrow^{n}}$. Following the transitions $B_{1},...,B_{i}$, we can apply the operator $\mathcal{T}_{F}$ to $\mathcal{T}_{F}{\uparrow^{n}}$. From Definition 2 and proposition 2, we know $c_{2}\in\mathcal{T}_{F}{\uparrow^{\alpha}}$ so a set of marked places $S\in\Omega$ must contain $c_{2}$. Therefore, $c_{2}$ is reachable from $c_{1}$ given an initial marking $m$.
@@
Theorem 4.1 says that a program transformed from an OEN is logically equivalent to the OEN based on the fixpoint semantics in Proposition 2 for evaluating OEN reachability queries. In the next section, we explain how we leverage boolean matrix operations to evaluate datalog transformed from OENs.
5 Boolean matrix logic programming
5.1 Problem setting
We propose the Boolean Matrix Logic Programming (BMLP) problem. In contrast to traditional logic program evaluation, BMLP uses boolean matrices to evaluate datalog programs.
Definition 3 (Boolean Matrix Logic Programming (BMLP) problem)
We focus on linear and immediately recursive datalog programs with at most two body literals to simulate a GEM. Our goal is to implement BMLP algorithms that can compute boolean matrices for evaluating large logic programs representing GEMs.
5.1.1 Creating boolean matrices
@@
We show boolean matrices created from $\mathcal{P}_{1}$ in Example 1. Individual substances are represented by constant symbols $c1$, $c2$ and $c3$. Sets of chemical substances are denoted by additional concatenated constant symbols, $c1\_c2$, $c3\_c4$. On the right-hand side, we create two boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ from the reaction clauses in $\mathcal{P}_{1}$. Since the first reaction_1 clause also contains the constant symbols of the other reaction_1 clauses, to avoid redundancy, only one binary code is created for these clauses in $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
$\mathcal{P}_{1}:\left\{\begin{aligned} ...&\\ reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ ...&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
@@
Constant symbols are represented by rows in $\textbf{R}_{1}$ and $\textbf{R}_{2}$. Constant symbols $c1\_c2$, $c3\_c4$ and $c5$ that appear on the left-hand side of reactions are denoted by binary bits in $\textbf{R}_{1}$. Constant symbols $c3$, $c4$, $c5$ and $c3\_c4$ that appear on the right-hand side of reactions are denoted by binary bits in $\textbf{R}_{2}$. Each row in the boolean matrices is written as a unit ground clause. For example, since $\mathcal{P}_{1}\models reaction(c1\_c2,c3\_c4)$, the first row of $\textbf{R}_{1}$ is written a binary code “00011” to represent $c1\_c2$ and the first row of $\textbf{R}_{2}$ is a binary code “01100” to represent $c3\_c4$. Each “1” binary bit indicates the presence of a constant symbol.
@@
We show the mapping of a single row matrix from the token definition from $\mathcal{P}_{1}$ in Example 1. Individual substances are represented by constant symbols $c1$ and $c2$. The set of chemical substances is denoted by an additional concatenated constant symbol, $c1\_c2$.
@@
The separation of this vector from $\textbf{R}_{1}$ and $\textbf{R}_{2}$ is meant to keep $\textbf{R}_{1}$ and $\textbf{R}_{2}$ static. Since the structure of the OEN does not vary often, we reduce the number of times that we need to re-create these two matrices while the definition of an initial marking usually changes across different simulations. We use all three created matrices, $\textbf{R}_{1}$, $\textbf{R}_{2}$ and v for bottom-up evaluation.
5.2 Iterative extension (BMLP-IE)
@@
Input: A $1\times n$ vector v, a $1\times k$ vector t that represents viable transitions, two $k\times n$ boolean matrices $\textbf{R}_{1}$, $\textbf{R}_{2}$ that encode unit ground clauses.
@@
Given a program $\mathcal{P}$ that contains a linear and immediately recursive clause $r$, we implement a BMLP algorithm called iterative extension to perform the bottom-up evaluation of $r$. The iterative extension algorithm (BMLP-IE) operates on two matrices to iteratively expand terms reachable in a partially grounded query $r(u,Y)$. Given that $\mathcal{P}$ contains $n$ constant symbols, we represent the constant symbols in $u$ as a $1\times n$ row vector or a $n$-bit binary code $\mathbf{v}$. Based on Section 5.1.1, we write $k\times n$ boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ for $k$ unit ground clauses used by $r$. $\textbf{R}_{1}$ and $\textbf{R}_{2}$ describe which constant symbols appear first and second in the unit ground clauses. In addition, a $1\times k$ boolean vector t is used to express additional control on the viability of transitions in the OEN. Transition can be turned “on” or “off” via this vector. “On” transitions have the binary value 1 and “off” transitions have the value 0. This allows us to temporarily modify the network connectivity without re-creating $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
@@
Theorem 5.1 shows that BMLP-IE has a polynomial runtime cubic in the size of the boolean matrices. Evaluating the program $\mathcal{P}$ using SLD resolutions has the same time complexity and the runtime of BMLP-IE primarily benefits from bit-wise operations. For SLD resolutions, the size of the subset of the Herbrand base provable from a query $r(u,Y)$ is bounded by $O(n)$. Each member of this subset can be found via a depth-first search. Since there are $k$ unit ground clauses and each variable can be grounded by at most $n$ constants, this query requires at most $n$ search in a tree of $O(k\times n)$ nodes. For OEN reachability queries, it is important to note that the size of boolean matrices grows linearly with the number of hypernodes because they are both bounded by the number of transitions.
6 Active learning
6.1 Hypothesis compression
We use the compression score of a hypothesis $h$ to optimise the posterior probability. A learner can derive the relative frequency $p(h)$ for $h$ to be chosen based on the encoding of $h$. In an optimal binary bit coding scheme shannon_mathematical_1963 , the length of $h$ denoted by $size(h)$ is $-log_{2}(p(h))$. The compactness of $h$’s encoding reflects its prior probability $p(h)=2^{-size(h)}$. We can use the compression function to compute the prior $p(h)$. Given a set of training examples $E$, the compression of a hypothesis progol ; ase_progol can be interpreted as
$compression(h,E)=size(E)-MDL_{h,E}$
$\frac{p(h|E)}{p(E|E)}=2^{compression(h,E)}$
We can re-arrange this to obtain the normalised posterior probability ase_progol :
The posterior probability is maximal when compression is maximal. By searching for a hypothesis with maximum posterior probability, a learner can maximise its expected predictive accuracy progol . To learn an accuracy hypothesis, the goal is to find a hypothesis with the highest compression score.
The compression function in $BMLP_{active}$ follows the Minimal Description Length (MDL) principle conklin_complexity-based_1994 . The most probable hypothesis $h$ for the training examples $E$ should be compact and have the fewest disagreements between $E$ and predictions made by $h$. This minimises $L(h)+L(E|h)$, where $L(h)$ is the descriptive complexity of $h$ and $L(E|h)$ is the descriptive complexity of encoding $E$ using $h$. We consider a compression function from ase_progol based on the MDL principle:
$compression(h,E)=|E^{+}|-\frac{|E^{+}|}{pc_{h}}(size(h)+fp_{h})$
where $E^{+}$ is the set of positive examples seen by the active learner, $pc_{h}$ is the positive coverage by $h$ and $fp_{h}$ is false positives covered by $h$. In other words, this compression function favours a general hypothesis with high coverage and penalises over-general hypotheses that incorrectly predict negative examples. The generality of $h$ can be estimated by computing its coverage $pc_{h}$ for a set of unlabelled instances progol . To determine the hypotheses’ compression scores, $BMLP_{active}$ looks at all available experiment instances to obtain a good estimate of a hypothesis’s generality. Learning with increasingly large hypothesis space and instance space thus requires efficient inferences in the background knowledge. Our BMLP approach addresses this by manipulating boolean matrices for bottom-up evaluations.
6.2 Expected cost of experiments
$EC(\emptyset,T)=0$
$EC({h},T)=0$
$EC(H,T)=min_{t\in T}\,[C_{t}+p(t)EC(H_{t},T-{t})+(1-p(t))EC(\overline{H_{t}},T-{t})]$
$+(1-p(t))(mean_{t^{\prime}\in T-\{t\}}C_{t^{\prime}})J_{\overline{H_{t}}}]$
where $p(t)$ is the probability that the outcome of the experiment $t$ is positive and $J_{H}=-\sum_{h\in H}p(h)\,log_{2}(p(h))$. The probability $p(h)$ is the prior of hypothesis $h$ calculated from the compression function. To estimate $p(t)$, we compute the sum of the probabilities of the hypotheses consistent with a positive outcome of $t$. While we only focus on auxotrophic mutant experiments in this work, users can flexibly define the experiment cost function for different types of experiments.
6.3 Active learning sample complexity
For some hypothesis space $H$ and background knowledge $BK$, let $V_{s}$ denote the version space of hypotheses consistent with $s$ training examples. For an active version space learner that selects one instance per iteration, $|V_{s}|$ denotes the size of the version space at the iteration $s$ of active learning. The shrinkage of the hypothesis space can be represented by the reduction ratio $\frac{|V_{s+1}|}{|V_{s}|}$ after querying the label of the $s$-th instance hocquette_how_2018 . The minimal reduction ratio $p(x_{s},V_{s})$ is the minority ratio of the version space $V_{s}$ partitioned by an instance $x_{s}$.
Definition 4
(Minimal reduction ratio hocquette_how_2018 ) The minimal reduction ratio over the version space $V_{s}$ by sampled instance $x_{s}$ is
$p(x_{s},V_{s})=\frac{min(\,|\{h\in V_{s}|\,h\cup BK\models x_{s}\}|,|\{h\in V_{s}|\,h\cup BK\not\models x_{s}\}|)}{|V_{s}|}$
The minimal reduction ratio of an actively selected instance can be computed before it is labelled. While in reality training examples might not be as discriminative, the optional selection strategy is to select instances with minimal reduction ratios as close as possible to $\frac{1}{2}$ with the ability to eliminate up to 50% of the hypothesis space. We describe the sample complexity advantage of active learning over random example selection by the following bound of an active version space learner’s sample complexity. A passive learner is considered a learner using random example selection since it does not have control over the training examples it uses.
Theorem 6.1 (Active learning sample complexity bound)

### v4

@@ Appendix B Equivalence between OEN and datalog
The transformation takes an OEN and an initial marking $m$. Let $t$ denote any transition. Input places of transitions are represented by $p^{*}$. A datalog program $\mathcal{P}$ is written with predicate symbols $t$, $r_{1}$ and $r_{2}$ and variables X, Y and Z:
  • write $token(m,p^{*})$ if the marking $m$ includes all $p^{*}$
@@
  • write $t(p_{1}^{*},p_{2}^{*})$ if $t$ connects $p_{1}^{*}$ and $p_{2}^{*}$
  • write $t(p^{*},q)$ if $t$ connects $p^{*}$ and a place $q$
@@
The transformed program is linear and immediately recursive since it contains only one recursive clause, and the recursive predicate appears only once in the head of this clause. We illustrate how reachable places from an initial marking can be computed by evaluating the transformed program.
@@
We use the following OEN to represent a metabolic pathway. Transitions $t_{1}$, $t_{2}$ and $t_{3}$ are described by $reaction\_1$, $reaction\_2$ and $reaction\_3$ clauses. We create two nodes $c1\_c2$ and $c3\_c4$. These transitions connect places $\{c_{3}$, $c_{4}$, $c_{5}\}$ which are represented by constants $\{c3$, $c4$, $c5\}$. For an initial marking $m_{1}=\{c_{1},c_{2}\}$, we can construct a datalog program $\mathcal{P}_{2}$:
$\mathcal{P}_{2}:\left\{\begin{aligned} token(m1,c1\_c2).&\\ token(m1,c1).&\\ token(m1,c2).&\\ reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ reaction(X,Y)\leftarrow token(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
The right-hand-side figures are the OEN represented by $\mathcal{P}_{2}$. The node $\{c_{3},c_{4}\}$ is reachable from $\{c_{1},c_{2}\}$ via the transition $t_{1}$. The place $c_{5}$ is reachable in two steps via transitions $t_{1}$ and $t_{2}$. We can compute these by evaluating groundings of $pathway(m1,Y)$ in the Herbrand model $M(\mathcal{P}_{2})$ which gives us $pathway(m1,c2)$, $pathway(m1,c3\_c4)$ and $pathway(m1,c5)$.
@@
It follows that all tokens at the places in the pre-set of a transition are consumed, and places in the post-set of a transition are assigned tokens.
@@
Let $(P,T,F)$ be an arbitrary OEN. Using Proposition 1, we show $\mathcal{T}_{F}$ over the lattice $(\mathcal{T}_{F},\subseteq)$ has a fixpoint. We refer to Tarski’s fixpoint theorem [68], which states a monotonic operator on a complete lattice has a least fixpoint. We prove that $\mathcal{T}_{F}$ is a monotonic operator.
@@
$\mathcal{T}_{F}$ is a monotonic function. According to Tarski’s fixpoint theorem [68], the function $\mathcal{T}_{F}$ has a least fixpoint $\mathcal{T}_{F}{\uparrow^{\alpha}}$ for some $\alpha\geq 0$. Since $\mathcal{T}_{F}$ is also a compact operator, if $p\in\mathcal{T}_{F}(\mathcal{T}_{F}{\uparrow^{\alpha}})$ then $p\in\mathcal{T}_{F}(\mathcal{T}_{F}{\uparrow^{n}})$ for some $n\geq 0$ and $\mathcal{T}_{F}{\uparrow^{n}}\subseteq\mathcal{T}_{F}{\uparrow^{\alpha}}$. The union of subsets of $\mathcal{T}_{F}{\uparrow^{\alpha}}$ is $\cup_{n\geq 0}\mathcal{T}_{F}{\uparrow^{n}}$, which covers every member of the least fixpoint. Therefore, $\mathcal{T}_{F}{\uparrow^{\alpha}}=\cup_{n\geq 0}\mathcal{T}_{F}{\uparrow^{n}}$.
@@
Prove the forward implication. Assume $\mathcal{P}\models r(c_{1},c_{2})$. Constant symbols $c_{1}$ and $c_{2}$ represent place nodes in the OEN. We know $r(c_{1},c_{2})$ is in the least model. Based on the definition of the Immediate Consequence Operator [71], $r(c_{1},c_{2})\leftarrow\,B_{1},...,B_{i}$ ($i\geq 0$) is a clause in $\mathcal{P}$ and $\{B_{1},...,B_{i}\}\subseteq\mathcal{T_{P}}{\uparrow^{\omega}}$. From Definition 3, we know that $B_{1},...,B_{i}$ are unit ground clauses describing the pre-sets and post-sets of transitions. There exists $n\geq 0$ such that $c_{1}\in\mathcal{T}_{F}{\uparrow^{n}}$. Following the transitions $B_{1},...,B_{i}$, we can apply the operator $\mathcal{T}_{F}$ to $\mathcal{T}_{F}{\uparrow^{n}}$. From Definition 4 and proposition 2, we know $c_{2}\in\mathcal{T}_{F}{\uparrow^{\alpha}}$ so a set of marked places $S\in\Omega$ must contain $c_{2}$. Therefore, $c_{2}$ is reachable from $c_{1}$ given an initial marking $m$.
@@
Theorem B.1 says that a program transformed from an OEN is logically equivalent to the OEN based on the fixpoint semantics in Proposition 2 for evaluating OEN reachability queries. In the next section, we explain how we leverage Boolean matrix operations to evaluate datalog transformed from OENs.
Appendix C Boolean matrix encoding
@@
We show Boolean matrices created from $\mathcal{P}_{2}$ in Example 1. Individual metabolites are represented by constant symbols $c1$, $c2$ and $c3$. Sets of chemical substances are denoted by additional concatenated constant symbols, $c1\_c2$, $c3\_c4$. On the right-hand side, we create two Boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ from the reaction clauses in $\mathcal{P}_{1}$. Since the first reaction_1 clause also contains the constant symbols of the other reaction_1 clauses, to avoid redundancy, only one binary code is created for these clauses in $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
$\mathcal{P}_{2}:\left\{\begin{aligned} ...&\\ reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ ...&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
@@
Constant symbols are represented by rows in $\textbf{R}_{1}$ and $\textbf{R}_{2}$. Constant symbols $c1\_c2$, $c3\_c4$ and $c5$ that appear on the left-hand side of reactions are denoted by binary bits in $\textbf{R}_{1}$. Constant symbols $c3$, $c4$, $c5$ and $c3\_c4$ that appear on the right-hand side of reactions are denoted by binary bits in $\textbf{R}_{2}$. Each row in the Boolean matrices is written as a unit ground clause. For example, since $\mathcal{P}_{2}\models reaction(c1\_c2,c3\_c4)$, the first row of $\textbf{R}_{1}$ is written a binary code “00011” to represent $c1\_c2$ and the first row of $\textbf{R}_{2}$ is a binary code “01100” to represent $c3\_c4$. Each “1” binary bit indicates the presence of a constant symbol.
@@
We show the mapping of a single row matrix from the token definition from $\mathcal{P}_{2}$ in Example 1. Individual metabolites are represented by constant symbols $c1$ and $c2$. The set of chemical substances is denoted by an additional concatenated constant symbol, $c1\_c2$.
@@
Appendix D Iterative extension (BMLP-IE)
@@
Input: A $1\times n$ vector v, a $1\times k$ vector t that represents transitions, two $k\times n$ Boolean matrices $\textbf{R}_{1}$, $\textbf{R}_{2}$ that encode unit ground clauses.
@@
The iterative extension algorithm (BMLP-IE) operates on two matrices to iteratively expand terms reachable in a partially grounded query $r(u,Y)$. Given that $\mathcal{P}$ contains $n$ constant symbols, we represent the constant symbols in $u$ as a $1\times n$ row vector or a $n$-bit binary code $\mathbf{v}$. We write $k\times n$ Boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ for $k$ unit ground clauses used by $r$. $\textbf{R}_{1}$ and $\textbf{R}_{2}$ describe which constant symbols appear first and second in the unit ground clauses. In addition, a $1\times k$ Boolean vector t is used to express additional control on the viability of transitions in the OEN. Transition can be turned “on” or “off” via this vector. “On” transitions have the binary value 1 and “off” transitions have the value 0. This allows us to temporarily modify the network connectivity without re-creating $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
@@
Theorem D.1 shows that BMLP-IE has a polynomial runtime cubic in the size of the Boolean matrices. Evaluating the program $\mathcal{P}$ using SLD resolutions has the same time complexity, and the runtime of BMLP-IE primarily benefits from bit-wise operations. For SLD resolutions, the size of the subset of the Herbrand base provable from a query $r(u,Y)$ is bounded by $O(n)$. Each member of this subset can be found via a depth-first search. Since there are $k$ unit ground clauses and each variable can be grounded by at most $n$ constants, this query requires at most $n$ search in a tree of $O(k\times n)$ nodes. For OEN reachability queries, it is important to note that the size of Boolean matrices grows linearly with the number of nodes because they are both bounded by the number of transitions.
Appendix E Active learning sample complexity
Theorem 4.1 shows the sample complexity gain of active learning over learning from randomly sampled instances.
Theorem E.1 (Active learning sample complexity bound)